# Exploiting Precision Mapping and Component-Specific Feature Enhancement for Breast Cancer Segmentation and Identification

### v4

@@ 3.1 Methodology
The ultrasound images are first augmented to handle class imbalance. Following augmentation, the images were preprocessed using a sequence of preprocessing steps – gamma correction, gaussian filtering, image resizing, and normalization. Pixel values in ultrasound images can reflect non-linearities, especially in high- or low-intensity regions. Gamma correction can help compensate for these non-linearities, leading to more accurate and visually appealing images. An effective technique for noise reduction and edge detail preservation in ultrasound images is the application of Gaussian filtering. This effectively reduces noise while preserving edge details. To preserve consistency throughout the dataset and facilitate batch processing, resizing is done to ensure the images fed into the proposed DL model have the same dimensions. Normalizing the image pixels to scale within a specific range enhances the quality of activation functions’ ability to capture the non-linearities in the data. Here, the images have been scaled to fall within the range (0, 1). The preprocessed images are then fed to the proposed Spatial-Channel Attention LinkNet Framework with InceptionResNet backbone for segmenting the tumor region. The segmented tumor maps are then fed to the proposed DCNNIMAF classifier to classify the segmented mass as benign, normal, or malignant. The overall workflow of this proposed work has been presented in Figure 1.
3.2 Dataset Exploration
@@
Figure 3: Training Data Distribution of Breast Ultrasound Images Before Augmentation
The rationale behind this augmentation approach is to level the data count of the ’normal’ and ’malignant’ classes, thereby aligning them more closely with the larger ’benign’ class. By increasing the training data for the ’normal’ and ’malignant’ classes through augmentation, the effects of class imbalance are aimed to be mitigated and enable the models to learn effectively from all classes. This approach ensures that the segmentation and classification models are trained on a more balanced dataset, thereby improving their ability to accurately segment, identify, and characterize breast tumors across different classes. This augmentation resulted in a well-balanced data distribution of each category, which has been represented in Figure 4.
Figure 4: Training Data Distribution of Breast Ultrasound Images After Augmentation
@@ 3.3 Preprocessing
Following augmentation, the images were preprocessed using a preprocessing pipeline, consisting of four stages – gamma correction, gaussian filtering, resizing, and image normalization. The output images obtained after each preprocessing step of breast ultrasound image preprocessing are shown in Figure 5.
@@ 3.3.1 Gamma Correction
Gamma correction serves as the initial preprocessing step tailored specifically for breast ultrasound images. It plays a pivotal role in enhancing the visibility of crucial anatomical structures and subtle details within the images. By adjusting the image’s brightness and contrast, gamma correction improves the delineation of tumor boundaries and enhances the visibility of tumor features. This step is particularly critical in breast cancer tumor segmentation, where accurate visualization of tumor margins is essential for precise delineation and subsequent analysis.
Gamma correction can be represented mathematically as follows:
@@ 3.3.2 Gaussian Filtering
Following gamma correction, Gaussian filtering is employed to mitigate speckle noise, a common artifact in ultrasound images that can obscure tumor boundaries and hinder accurate segmentation. By selectively smoothing out noise while preserving essential details, Gaussian filtering improves the clarity of tumor features and enhances the accuracy of segmentation algorithms. This step is crucial in breast cancer tumor segmentation and classification, as it reduces noise artifacts and improves the fidelity of tumor delineation, leading to more accurate and reliable segmentation results. Gaussian filtering ensures that the images are cleaner and more conducive to subsequent segmentation and classification tasks, facilitating the accurate identification and characterization of breast tumors.
Gaussian filtering is given by:
@@ 3.3.3 Ultrasound Image Resizing
Once the images have undergone gamma correction and Gaussian filtering, resizing is performed to standardize image dimensions, facilitating compatibility with segmentation and classification algorithms. Standardized image dimensions are essential for ensuring consistency and comparability across different datasets and analysis pipelines. Resizing enables researchers to create a uniform framework for analysis, simplifying the processing pipeline and reducing computational complexity.
Resizing of breast cancer images can be represented mathematically as:
@@ 3.3.4 Pixel Normalization
The last preprocessing stage, normalization, scales the pixel values of images to a standardized range, typically from 0 to 1. This normalization process is crucial for ensuring consistency in pixel intensity across different images, which is essential for training machine learning models and neural networks. Normalization enhances the comparability of images and improves the convergence speed of machine learning algorithms during training. By eliminating variations in intensity that may arise due to differences in acquisition parameters or imaging conditions, normalization ensures that segmentation and classification algorithms can learn effectively from the data, leading to more accurate and reliable analysis results.
The normalization process is denoted as:
@@
3.4 Dual Attention and CNN Backbone Enhanced LinkNet Framework for Breast Cancer Segmentation
This section presents the proposed framework for breast cancer tumor segmentation utilizing a LinkNet framework with an InceptionResNet backbone, employing a dual spatial-channel attention mechanism. The framework takes preprocessing breast ultrasound images and their corresponding ground truth masks as input to the segmentation model and provides the predicted segmentation map as output.
The LinkNet architecture [29] is a deep learning model designed for semantic segmentation tasks, particularly in the context of biomedical imaging. The encoder of the proposed framework is built using the InceptionResNet CNN model [30], which is designed to capture contextual information from the entire input image. The decoder is a series of transpose convolution layers with dual spatial-channel attention mechanisms incorporated within the decoder blocks.
3.4.1 Encoder Section
The encoder section of the segmentation architecture is designed using an InceptionResNet CNN backbone and thus consists of a stem block, three types of InceptionResNet blocks, and two types of reduction blocks.
The stem block begins with three convolution layers and is followed by a max pooling layer and a convolution layer where the layers get executed at the same time. This is followed by a filter concatenation layer and this is split into two paths that are parallel to each other. One of the paths contains two convolutional layers while the other is composed of four convolutional layers. Both paths are combined using filter concatenation and are followed by a parallel convolution and max pooling again which is further followed by filter concatenation.
@@ 3.4.1 Encoder Section
3.4.2 Decoder Section
The Decoder section is composed of decoder blocks, spatial-channel attention blocks, convolution and transpose convolution layers, and a softmax activation function. The decoder block begins with convolution and a batch normalization operation followed by a transpose convolution operation and another batch normalization operation which is followed by convolution and batch normalization operations again.
@@ 3.4.2 Decoder Section
The spatial channel attention block begins with two double convolution operations taking place simultaneously followed by the addition of the two feature maps obtained from the operation. The addition operation is followed by the introduction of non-linearity using ReLU activation followed by another convolution operation. The convolution operation is followed by a sigmoid activation function to restrict the values to lie within the range 0 and 1.
@@
This is followed by channel attention. The Channel attention block is composed of two different pooling operations (max pooling and average pooling) which happen simultaneously and the obtained feature maps are given as input to a shared multi-layered perceptron. The shared MLP is composed of a flatten layer, gaussian error linear unit (GELU) activation function and dropout layers. Once these three operations are done flatten and dropout operations are performed again. The decoder operation ends with a transpose convolution operation, two convolution operations, and a softmax activation function thus displaying the segmented output.
@@
Figure 6: Spatial-Channel Attention LinkNet Framework with InceptionResNet Backbone Layer Architecture for Breast Cancer Segmentation
3.4.3 Workflow and Execution
Initially, the image is processed through a stem block, which captures low-level features such as edges and textures. These edges outline the boundaries of potential tumors, while textures reveal the internal structure of these masses, which often differ significantly between healthy tissue and malignancies. Following the stem block, the image progresses through five InceptionResNet-A Blocks. Microcalcifications require finer resolution, whereas architectural distortions span larger areas. The residual connections, facilitate deeper networks by mitigating the vanishing gradient problem, and capture a broad spectrum of features at various scales. This multi-scale feature capture is crucial for analyzing ultrasound images of breast tissue, where abnormalities can manifest at different scales.
Subsequently, the image encounters a reduction block, which reduces the spatial dimensions of the feature maps. This reduction allows the model to focus on higher-level features and significantly reduces computational complexity, facilitating more efficient processing. This is particularly useful for identifying broader patterns indicative of cancer, such as the overall shape and orientation of a mass.
The image then navigates through two InceptionResNet Block B layers, further through another stem block. This refines the detection of mid-level features, such as more nuanced textural patterns and subtle edge variations. The stem block repetition extracts additional low-level features that complement the more complex features identified in the intermediate stages, ensuring the model has a comprehensive grasp of the ultrasound image’s content.
Following this, the image passes through a reduction block, which reduces the spatial dimensions of the feature maps. This reduction allows the model to focus on low-level features, essential for the precise delineation of tumor boundaries. The image then enters five InceptionResNet Block C layers, and finally into an average pooling layer. These operations are optimized for extracting high-level semantic features for differentiating between various types of tissues present in the image.
This compressed feature map is then fed to the LinkNet decoder which transforms the abstracted feature map into a spatially coherent segmentation map. In the decoder, upsampling refines the segmentation map generated by the encoder, and the attention mechanism focuses specifically on the tumor region, enhancing its emphasis. By integrating spatial and channel attention mechanisms, the model can enhance feature maps by emphasizing spatial locations and informative channels. This comprehensive approach improves the model’s capability to understand intricate tumor patterns and structures, thereby enhancing segmentation performance.
Initially, the feature map is fed to 2 convolutional blocks, followed by a spatial-channel attention block, which is repeated thrice. They perform a preliminary enhancement of the map, focusing on sharpening the details and adjusting the contrast to make the underlying structures more prominent. This ensures that the feature map contains clear and distinguishable elements that correspond to the anatomical structures within the breast ultrasound images. It is then passed to the first decoder block.
The initial decoder block is designed to capture high-level semantic features essential for segmenting larger structures within breast ultrasound images. It facilitates the reconstruction of the spatial relationships and contextual information abstracted away during the encoding process. The spatial-channel attention block that follows this decoder block scrutinizes the feature map to identify and accentuate the regions that are most likely to contain tumor structures. This is achieved by assigning higher weights to the spatial locations that exhibit characteristics typical of tumors, such as irregular shapes and unusual textural patterns. The channel attention mechanism analyses the feature map across different channels to determine the ones that carry the most relevant information for segmentation. By amplifying the signals from these informative channels, the model can better discern the unique features that differentiate tumor tissue from the surrounding healthy tissue.
Finer textures and structures within the tumors are captured as the feature map moves up to the second decoder block. The spatial-channel attention block adjusts the feature map’s weights to emphasize the spatial locations where these detailed features are most prominent, resulting in more precise segmentation of smaller tumor components. Channel attention further identifies the most relevant feature map channels for the task, focusing on the texture and shape of the tumors.
In the third decoder block, the feature map captures even more detailed features, including intricate patterns and structures within the tumors. The attention mechanism in this block focuses on the boundaries of the tumor region, which helps the model improve the quality of the produced segmentation map. This makes the output more accurate and minimizes extraneous markings.
The final decoder block is responsible for capturing the most detailed features, including the specific patterns and structures unique to each tumor. The attention mechanism allows the model to distinguish between benign and malignant types of tumors and identify subtle variations within a single tumor type. The output from this decoder block is transpose-convolved to ensure a consistent output shape of the segmentation map, followed by convolutions to correct the output channels. This finally transforms the abstracted feature map into a detailed and accurate segmentation map.
@@ 3.4.3 Workflow and Execution
The model was trained by backpropagating over a custom loss function (21), equal to an aggregate of focal loss (19) and dice (Jaccard) loss (20) obtained after each training epoch.
@@
The model specifications and parameters of the proposed Spatial-Channel Attention LinkNet Framework with InceptionResNet Backbone are shown in Table 1.
Table 1: Dual Attention and CNN Backbone Enhanced LinkNet Segmentation Framework
@@
3.5 Multi-Attention Integrated Deep CNN Framework for Breast Cancer Classification
This section presents the proposed breast cancer deep learning classification model, coined Deep CNN with an Integrated Multi-Attention Framework (DCNNIMAF). Utilizing multiple attention modules integrated within its architecture the proposed approach is designed to effectively classify breast ultrasound images into malignant, benign, or normal categories. The input to the model comprises preprocessed breast ultrasound images and outputs the predicted class to which the image belongs.
@@ 3.5 Multi-Attention Integrated Deep CNN Framework for Breast Cancer Classification
The model architecture of DCNNIMAF integrates several pivotal blocks designed to extract pertinent features from the input breast ultrasound images. These blocks include convolutional blocks, double convolutional blocks, self-attention blocks, and fully connected layers. Each block plays a crucial role in feature extraction and classification. The layer architecture diagram of the proposed DCNNIMAF model is shown in Figure 7.
Figure 7: DCNNIMAF Classification Layer Architecture for Breast Cancer Classification
3.5.1 Convolutional Block
The convolutional block within DCNNIMAF consists of a convolutional layer, followed by a batch normalization layer, and finally an activation layer. The activation function used varies between Leaky ReLU and SiLU in different convolutional blocks. The operations performed by the block on the input feature map are mathematically represented as follows:
$\underbrace{O_{i,j}}_{\text{Output feature map value}}=\sum_{m=0}^{M-1}\sum_{n=0}^{N-1}\underbrace{I_{i+m,j+n}}_{\text{Input pixel value}}\cdot\underbrace{F_{m,n}}_{\text{Filter weight}}+\underbrace{b}_{\text{Bias}}$ (22)
$\underbrace{\text{BN}(x)}_{\text{Batch Normalization}}=\underbrace{\gamma}_{\text{Scale parameter}}\frac{\underbrace{x}_{\text{Input value}}-\underbrace{\mu}_{\text{Mean}}}{\sqrt{\underbrace{\sigma^{2}}_{\text{Variance}}+\underbrace{\epsilon}_{\text{Small constant}}}}+\underbrace{\beta}_{\text{Shift parameter}}$ (23)

### v5

@@ 3.1 Methodology
Augmentation is applied initially to the ultrasound images dealing with class imbalance. In the subsequent stage, a series of preprocessing steps comprising gamma correction, gaussian filtering, resizing, and normalizing are used on these images. Pixel values are known to contain non-linearity in ultrasound images primarily in the regions of very high or low intensity. Gamma correction would be apt to counter this effect, therefore producing images more accurate and appealing to the human eye. An effective technique for noise reduction and edge detail preservation in ultrasound images is the application of Gaussian filtering. This effectively reduces noise while preserving edge details. To preserve consistency throughout the dataset and facilitate batch processing, resizing is done to ensure the images fed into the proposed DL model have the same dimensions. Normalizing the image pixels to scale within a specific range enhances the quality of activation functions’ ability to capture the non-linearities in the data. Here, the images have been scaled to fall within the range (0, 1). The preprocessed images are then fed to the proposed Spatial-Channel Attention LinkNet Framework with InceptionResNet backbone for segmenting the tumor region. The segmented tumor maps are then fed to the proposed CSFEC-Net classifier to classify the segmented mass as benign, normal, or malignant. The overall workflow of this proposed work has been presented in Figure 1.
3.2 Dataset Description
@@
Figure 3: Training and Validation Dice Coefficient Curves of Proposed Localization Framework
Figure 4: Training and Validation IoU Score Curves of Proposed Localization Framework
The rationale behind this augmentation approach is to level the data count of the ’normal’ and ’malignant’ classes, thereby aligning them more closely with the larger ’benign’ class. By augmenting the training data for the ’normal’ and ’malignant’ classes, it is hoped that the effects of class imbalance are to be mitigated and enable the models to learn effectively from all classes. This approach ensures that the segmentation and classification models are trained on a more balanced dataset, thus improving their ability to accurately segment, identify, and characterize breast tumors across different classes. This augmentation resulted in a well-balanced data distribution of each category, which has been represented in Figure 4.
@@ 3.3 Preprocessing
Following augmentation, the images were preprocessed using a pipeline, composed of four stages – gamma correction, gaussian filtering, resizing, and image normalization. The output images obtained after each preprocessing step of breast ultrasound image preprocessing are shown in Figure 5.
@@ 3.3.1 Gamma Correction
Gamma correction is the first preprocessing step in our ultrasound image preprocessing pipeline. It has a very important role in enhancing the visibility of key anatomical structures and subtle details in the images. Gamma correction enhances the boundary delineation of tumors and improves the visibility of tumor features by adjusting the brightness and contrast of the image. This is especially crucial for the step in breast cancer tumor segmentation since correct visualization of tumor margins helps to properly define the tumors and the regions around them for better analysis. In this, gamma correction maps the input pixel intensity, $I_{\text{in}}$, into the output pixel intensity, $I_{\text{out}}$, as determined by the value of the gamma $\gamma$ given in Equation 1.
@@ 3.3.2 Gaussian Filtering
After gamma correction, we apply Gaussian filtering in order to reduce speckle noise, a prevalent artefact in ultrasound images which can obscure the boundaries of a tumor and prevent its proper segmentation. By selectively smoothing out noise while preserving necessary details, Gaussian filtering improves the clarity of tumor features and enhances the accuracy of our segmentation algorithms. This step is critical for segmenting and classifying tumor in breast cancer: since it reduces noise artifacts as well as increases the fidelity of tumor delineation to the extent that more accurate and reliable segmentation results are obtained. In summary, Gaussian filtering ensures that our images are cleaner and, by this, more amenable to subsequent segmentation and classification steps, thus allowing us to detect and characterize the tumors properly. The mathematical representation of the Gaussian filtering process is as follows (Equation 2), where $I_{\text{in}}(x,y)$ represents the input pixel intensity at coordinates $(x,y)$, $I_{\text{out}}(x,y)$ represents the output pixel intensity after Gaussian filtering, and $G(i,j)$ represents the Gaussian kernel applied to the neighborhood of the pixel. The size of the Gaussian kernel is represented by $N$.
@@ 3.3.3 Ultrasound Image Resizing
[…]
@@ 3.3.4 Pixel Normalization
After gamma correction and then application of the Gaussian filtering to the images, normalization will be used to map pixel values within the normalized image to an appropriate, consistent range - usually to lie between 0 and 1. This procedure normalizes the intensity of in-input pixels to ensure equivalence between different images as required when preparing images to train neural networks. By transforming the input image $I_{\text{in}}$ to the output image $I_{\text{out}}$, where
@@
we eliminate variations in intensity that may arise due to differences in acquisition parameters or imaging conditions. The algorithmic workflow of our preprocessing pipeline is presented in Algorithm 1.
Algorithm 1 Breast Ultrasound Image Preprocessing
0:  2D breast ultrasound images
0:  Preprocessed breast ultrasound images
1:  Function preprocess_image():
[…]
4:  for $h\leftarrow 0$ to $height$ do
5:     for $w\leftarrow 0$ to $width$ do
6:        $gamma\_corrected\_image(h,w)\leftarrow\text{input}(h,w)^{\gamma}$
7:     end for
8:  end for
9:  for $h\leftarrow 0$ to $height$ do
10:     for $w\leftarrow 0$ to $width$ do
11:        $gaussian\_filtered\_image(h,w)\leftarrow$
12:                  $\sum_{i=-h/2}^{h/2}\sum_{j=-w/2}^{w/2}gamma\_corrected\_image(h+i,w+j)\cdot G(i,j)$
13:        $G(i,j)=\dfrac{1}{2\pi\sigma^{2}}\exp\left(-\dfrac{i^{2}+j^{2}}{2\sigma^{2}}\right)$
14:     end for
15:  end for
16:  $r_{h},r_{w}\leftarrow\text{resizing factors for image of size }(height,width)$
17:  $h^{\prime},w^{\prime}\leftarrow\dfrac{height}{r_{h}},\ \dfrac{width}{r_{w}}$
18:  $resized\_image\leftarrow gaussian\_filtered\_image[r_{h}{:}h^{\prime},\ r_{w}{:}w^{\prime}]$
19:  for $h\leftarrow 0$ to $h^{\prime}$ do
20:     for $w\leftarrow 0$ to $w^{\prime}$ do
21:        $normalised\_image(h,w)\leftarrow\dfrac{resized\_image(h,w)-\min(resized\_image)}{\max(resized\_image)-\min(resized\_image)}$
22:     end for
23:  end for
24:  $preprocessed\_image\leftarrow normalised\_image$
25:  return  $preprocessed\_image$
3.4 Precision Mapping and Attention-Driven LinkNet (PMAD-LinkNet) Framework for Breast Lesion Segmentation
This section presents the proposed PMAD-LinkNet framework for breast cancer tumor segmentation. It takes preprocessing breast ultrasound images and their corresponding ground truth masks as input and produces the predicted segmentation map as output.
The LinkNet architecture [29] is a neural network model, used for semantic segmentation especially in the field of biomedical imaging. The encoder of the proposed framework is built using an InceptionResNet backbone [30] which aims to take into consideration the context of the whole image input. The decoder is composed of the series of the transpose convolution layers that have PMM modules and dual-attentions incorporated within the decoder blocks. The layer architecture of the proposed PMAD-LinkNet is presented in Figure 6.
Figure 6: PMAD-LinkNet Layer Architecture for Breast Cancer Segmentation
3.4.1 PMAD-LinkNet Encoder
The encoder section of the segmentation architecture consists of a stem block, three types of InceptionResNet blocks, and two types of reduction blocks.
The stem block starts off with a series of three convolutions before proceeding to a max pooling layer and a convolution layer which occur in parallel. After that comes a filter concatenation layer which in turn branches out into two parallel paths. One of the branches has two convolutional layers while the other branch has four of them. Both branches rejoin through filter concatenation and are in turn followed again by a parallel convolution and max pooling and then another filter concatenation step.
@@ 3.4.1 PMAD-LinkNet Encoder
3.4.2 PMAD-LinkNet Decoder
The decoder section of the PMAD-LinkNet segmentation framework integrates the PMM module within its decoder blocks to achieve precise pixel-level refinement of tumor boundaries. Each decoder block comprises convolutional layers, batch normalization, transpose convolution layers, spatial-channel attention blocks, and the PMM module, culminating with a softmax activation function for the final segmentation output.
The decoder block begins with a convolution operation followed by batch normalization, as described in Equation (10). Batch normalization stabilizes the learning process by normalizing the input features using the mean and variance computed over the mini-batch:
@@ 3.4.2 PMAD-LinkNet Decoder
A transpose convolution operation follows, which upsamples the feature maps to recover spatial resolution lost during encoding. This operation is mathematically represented in Equation (11):
@@
After another batch normalization, the Precision Mapping Mechanism (PMM) is introduced. The PMM dynamically calibrates spatial mappings through a spatial-channel attention block by analyzing morphological variations within the feature maps. It computes adaptive spatial adjustments for each pixel, effectively refining the feature map to accurately localize tumor edges. This is achieved through a learned refinement process that compensates for inconsistencies in image quality and variations in tumor morphology. It ensures that the spatial mappings are precisely aligned with the tumor boundaries at the pixel level. It begins with two parallel double convolution operations, capturing diverse spatial features. The resulting feature maps are then combined using an addition operation as per Equation (12). A ReLU activation introduces non-linearity, followed by another convolution operation. The output is passed through a sigmoid activation function (Equation (13)) to normalize the feature values between 0 and 1.
@@
The pooled features are input to a shared multi-layer perceptron (MLP) consisting of flatten layers, Gaussian Error Linear Unit (GELU) activation functions, and dropout layers. The GELU activation function is defined in Equations (15) and (16). Dropout regularization is applied to prevent overfitting, as shown in Equation (17).
@@
The PMM reweights the feature maps, emphasizing the most informative features for segmentation. The decoder operation concludes with a transpose convolution, two convolution operations, and a softmax activation function  (18) to generate the segmented output.
@@
By integrating the PMM within the decoder blocks, the PMAD-LinkNet framework dynamically adapts spatial mappings based on morphological variations. The PMM’s adaptive mapping strategy provides pixel-level precision in localizing tumor edges by adjusting spatial mappings through learned refinements. This approach effectively compensates for inconsistencies in image quality and variations in tumor morphology, leading to enhanced segmentation accuracy and precise delineation of tumor boundaries. The segmentation process is algorithmically presented in Algorithm 2.
Algorithm 2 PMAD-LinkNet for Breast Lesion Segmentation
[…]
0:  Trained segmentation framework $\text{model}_{\text{segmentation}}$
1:
[…]
3:  $\beta\leftarrow$ batch size
4:  $T\leftarrow$ total epochs
5:  $\theta\leftarrow$ parameters of $\text{model}_{\text{segmentation}}$
6:  $n\leftarrow$ total classes
7:  Define loss functions:
8:   $\mathcal{L}_{1}\leftarrow$ Jaccard loss
9:   $\mathcal{L}_{2}\leftarrow$ Dice loss
10:   $\mathcal{L}_{3}\leftarrow$ Categorical Focal loss
11:  for epoch $\leftarrow 1$ to $T$ do
12:     $\mu\leftarrow$ learning rate
13:     for each batch $(x,\text{mask}_{\text{GT}})$ sampled from training data do
14:        $\hat{y}\leftarrow\text{model}_{\text{segmentation}}(x)$
15:        $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{1}(\hat{y},\text{mask}_{\text{GT}})+\mathcal{L}_{2}(\hat{y},\text{mask}_{\text{GT}})+\mathcal{L}_{3}(\hat{y},\text{mask}_{\text{GT}})$
16:        Update $\theta$ using backpropagation with loss $\mathcal{L}_{\text{total}}$:
17:        $\theta\leftarrow\theta-\mu\nabla_{\theta}\mathcal{L}_{\text{total}}$where $\nabla_{\theta}\mathcal{L}_{\text{total}}$ is the gradient of the loss with respect to $\theta$}
18:     end for
19:     if validation performance plateaus then
20:        Adjust learning rate $\mu$ to promote further learning
21:     end if
22:  end for
23:  $\text{output}_{\text{segmentation}}\leftarrow\text{model}_{\text{segmentation}}$
24:  return  $\text{output}_{\text{segmentation}}$
3.4.3 Training Setup and Specifications
@@
The model was trained by backpropagating over a custom loss function (Equation 21), equal to an aggregate of focal loss (Equation 19) and dice (Jaccard) loss (Equation 20) obtained after each training epoch.
@@
The model specifications and parameters of the proposed PMAD-LinkNet segmentation framework are shown in Table 1.
Table 1: PMAD-LinkNet Segmentation Framework Specifications
@@
3.5 Component-Specific Feature-Enhanced Classifier (CSFEC-Net) for Breast Cancer Classification
This section presents the proposed Component-Specific Feature-Enhanced Classifier (CSFEC-Net) for breast lesion classification. The input to the model comprises segmented tumors and outputs the predicted class to which the lesion belongs.
@@ 3.5 Component-Specific Feature-Enhanced Classifier (CSFEC-Net) for Breast Cancer Classification
The model architecture of CSFEC-Net integrates several pivotal blocks designed to extract pertinent features from the input breast ultrasound images. These blocks include convolutional blocks, double convolutional blocks, self-attention blocks, and fully connected layers. Each block plays a crucial role in feature extraction and classification. The layer architecture diagram of the proposed CSFEC-Net model is shown in Figure 7.
Figure 7: CSFEC-Net Classification Layer Architecture for Breast Cancer Classification
3.5.1 Workflow and Execution
The classification module of the Component-Specific Feature-Enhanced Classifier (CSFEC-Net) is meticulously designed to accurately differentiate between malignant, benign, and normal breast tissues. Central to this module is the novel Component-Specific Feature Enhancement Module (CSFEM), which employs dynamic multi-attention mechanisms to selectively amplify discriminative features across different tissue types. The information flow within the CSFEC-Net commences with an input layer of dimensions $256\times 256\times 3$, corresponding to the input segmented tumor.
Initially, the segmentation map is processed through a convolutional block (5) comprising 512 filters, a padding of 2, and a kernel size of 3. This block extracts low-level features such as textures and edges from the input image. The output from the first convolutional block is then passed through a second convolutional block with 256 filters, identical kernel size, and padding, but incorporates the Sigmoid Linear Unit (SiLU) activation function (22) to introduce non-linearity for enhanced feature discrimination.
A double convolutional block is then employed, consisting of two consecutive convolutional layers with 256 filters each. This sequential processing facilitates the extraction of deeper and more abstract features, which are crucial for distinguishing between different breast tissue characteristics indicative of cancerous growth. The integration of the Component-Specific Feature Enhancement Module (CSFEM) follows, dynamically enhancing discriminative features specific to benign, malignant, and normal tissues. The CSFEM employs a multi-level attention approach, leveraging both spatial and self attention mechanisms to magnify distinguishing features. The attention-enhanced feature map $F^{\prime}$ is computed as:
$F^{\prime}=F\otimes\text{SA}(F)\otimes\text{CA}(F)$ (23)
where $F$ represents the input feature map, $\text{SA}(F)$ denotes the spatial attention map, $\text{CA}(F)$ denotes the channel attention map and $\otimes$ signifies element-wise multiplication. To capture features at multiple scales, the CSFEM incorporates a multi-scale attention pipeline:
$\text{MSA}(F)=\text{Concat}(\text{Conv}(F,1),\text{Conv}(F,3),\text{Conv}(F,5))\cdot W$ (24)
where $\text{Conv}(F,k)$ denotes a convolution operation with kernel size $k$, and $W$ represents the learnable weights for feature aggregation. The enhanced features are obtained by applying the multi-scale attention maps to the original feature maps:
$F_{\text{enhanced}}=F+\text{MSA}(F)\otimes F$ (25)
This process increases inter-class separability and intra-class consistency by selectively amplifying discriminative traits of different tissue types, effectively addressing overlapping feature spaces in heterogeneous tumor segmentations. It enhances the model’s ability to focus on relevant cellular distributions and the presence of necrosis, filtering out irrelevant information and potential artifacts that could obscure diagnosis.
Following the spatial attention in the CSFEM, the feature maps undergo additional convolutional blocks to further distill discriminative representations. A convolutional block with 128 filters, a kernel size of 4, and padding of 2, coupled with a Leaky ReLU activation function (26), refine the features. Subsequently, another convolutional block with 128 filters, padding of 1, and a kernel size of 3 using SiLU activation is applied. Two additional convolutional blocks with 128 and 64 filters, kernel sizes of 4 and 3, and appropriate padding are utilized before feeding the features into a spatial attention mechanism. This mechanism adjusts the feature maps to accentuate subtle differences between various tissue characteristics associated with malignant and benign tumors.